OpenStudy (anonymous):

The number of mosquitoes in millions in a certain area, M(x), depends on the June rainfall in inches, x, approximately as follows M(x)=10x-x^2 1. find the rainfall that will produce the maximum number of mosquitoes. 2.What is the maximum number of mosquitoes?

OpenStudy (bahrom7893):

first one is asking for what x will m(x) be maximized

OpenStudy (bahrom7893):

m' = 10-2x = 0 x = 5

OpenStudy (bahrom7893):

plug that back into the original: M(5) = 50-25 = 25